# From multitype branching Brownian motions to branching Markov additive processes

### v1

@@ 3.1 Truncated derivative martingales
For $x\in\mathbb{R}$, let $\tau_{x}:=\inf\{t\geq 0\colon\widehat{X}_{\xi}(t)<x\}$. According to [DLK17, Theorem 29], there exists finite and non-negative functions $(R_{i})_{i\in\mathcal{I}}$, such that the process
@@
Since $\widetilde{\mathbf{E}}_{x,i}^{\theta^{*}}[(-\widehat{X}_{\xi})^{2}(t)]<\infty$, we know from [DLK17, Theorem 35 and Lemma 38] that $(-\widehat{X}_{\xi},J_{\xi})$ has tight overshoot under $\widetilde{\mathbf{P}}_{x,i}^{\theta^{*}}$, and equivalently, $\sum_{j\in\mathcal{I}}\widetilde{\pi}_{j}\widetilde{\mathbf{E}}_{0,i}^{\theta^{*}}[H^{-}(1)]\in(0,\infty)$, where $H^{-}$ is the so-called descending ladder height process of $(\widehat{X}_{\xi},J_{\xi})$ under $\widetilde{\mathbf{P}}_{0,i}^{\theta^{*}}$ and $\widetilde{\pi}$ is the stationary distribution of $J_{\xi}(t)$ under $\widetilde{\mathbf{P}}_{0,i}^{\theta^{*}}$.  [^6] Note that, as remarked by [DLK17, Page 1995], we do not require non-lattice assumption. Then it follows from the Markov renewal theory (see e.g. [Lalley84] or [DLK17, Theorem 28 (i)]; note that the version we use here holds for the lattice case as well) that
@@
$\mathbf{E}_{x,i}\left[Z_{\theta^{*}}^{(b)}(t)\right]$ $=V_{i}(\theta^{*})e^{-\theta^{*}x}\widetilde{\mathbf{E}}_{x,i}^{\theta^{*}}\left[R_{J_{\xi}(t)}(\widehat{X}_{\xi}(t)+b)\mathbbm{1}_{\left\{\inf_{s\leq t}\widehat{X}_{\xi}(s)\geq-b\right\}}\right]=V_{i}(\theta^{*})e^{-\theta^{*}x}R_{i}(\theta^{*}x+b).$
@@
The convergence of the derivative martingale, stated in Theorem 1.6, would rely on the $\mathcal{L}^{1}$ convergence of the truncated derivative martingale $Z_{\theta^{*}}^{(b)}$. We give the following theorem.
@@
Then $\mathbb{P}_{x,i}^{\uparrow}$ is known as the law of a MAP conditioned to stay positive; we refer to [DLK17, Appendices A.7-A.8] for more details. Lévy processes (with single type) conditioned to stay positive have been the subject of a large literature; we refer to [ChD05] and reference therein. Define $\mathbf{P}_{x,i}^{\uparrow}$ as the projection of $\widetilde{\mathbf{P}}_{x,i}^{\uparrow}$ on $(\mathcal{T},\mathcal{F})$. By projection (2.24) and change of measure (2.18), we also deduce the connection with the original law $\mathbf{P}_{x,i}$: for $t\geq 0$,
@@
Denote by $\mathbf{M}(\mathrm{d}s,\mathrm{d}k)$ the counting measure (on $\mathbb{R}_{+}\times\mathbb{Z}_{>0}$) of the spine’s fission times and number of children. Let $\mathcal{G}$ be the $\sigma$-field generated by the spine $(X_{\xi},J_{\xi})$ and $\mathbf{M}$, as given in Lemma 2.43. By [ShiBRW, Lemma 4.2], to show that $(Z_{\theta^{*}}^{(b)}(t),t\geq 0)$ is uniformly integrable, and $\mathbf{E}_{x,i}\left[\frac{Z_{\theta^{*}}^{(b)}(\infty)}{Z_{\theta^{*}}^{(b)}(0)}\right]=1$, it suffices to prove that
@@
$\widetilde{\mathbf{E}}_{x,i}^{\uparrow}\left[Z_{\theta^{*}}^{(b)}(t)\,\middle|\,\mathcal{G}\right]=R(b+X_{\xi}(t))e^{-X_{\xi}(t)}+\int_{[0,t]\times\mathbb{Z}_{>0}}kR_{J_{\xi}(s)}(b+X_{\xi}(s))V_{J_{\xi}(s)}(\theta^{*})e^{-X_{\xi}(s)}\mathbf{M}(\mathrm{d}s,\mathrm{d}k).$
As $X_{\xi}(t)$ under $\widetilde{\mathbf{P}}^{\uparrow}_{x,i}$ is a MAP conditioned to stay above b-b, by [DLK17, Proposition 33] we have $\lim_{t\to\infty}X_{\xi}(t)=+\infty$ $\widetilde{\mathbf{P}}^{\uparrow}_{x,i}$-a.s. Therefore the first term above converges to 0, $\widetilde{\mathbf{P}}^{\uparrow}_{x,i}$-a.s.
@@
Recall that $\mathbf{M}$ is a Cox process, in the sense that given $(X_{\xi}(t),{J}_{\xi}(t))_{t\geq 0}$, the conditional distribution of $\mathbf{M}$ is a Poisson point process with intensity $\widetilde{\beta}_{{J}_{\xi}(t)}\mathrm{d}t\otimes\mathrm{d}\widetilde{\mu}_{{J}_{\xi}(t)}$, with each $\widetilde{\mu}_{i}$ the size-biased offspring distribution. Let $C_{\mathcal{I}}:=\max_{i\in\mathcal{I}}\widetilde{\beta}_{i}>0$. The compensation formula leads to
$\widetilde{\mathbf{E}}^{\uparrow}_{x,i}[A_{1}]$ $\leq C_{\mathcal{I}}\widetilde{\mathbf{E}}^{\uparrow}_{x,i}\bigg[\int_{\mathbb{R}_{+}}R_{J_{\xi}(s)}(b+X_{\xi}(s))V_{J_{\xi}(s)}(\theta^{*})e^{-X_{\xi}(s)}\Big(\sum_{k\geq 1}\widetilde{\mu}_{J_{s}}(k)k\mathbbm{1}_{\left\{k\leq e^{\frac{1}{3}X_{\xi}(s)}\right\}}\Big)\mathrm{d}s\bigg]$
$\leq C_{\mathcal{I}}\int_{\mathbb{R}_{+}}\widetilde{\mathbf{E}}^{\uparrow}_{x,i}\left[R_{J_{\xi}(s)}(b+X_{\xi}(s))V_{J_{\xi}(s)}(\theta^{*})e^{-\frac{2}{3}X_{\xi}(s)}\right]\mathrm{d}s$
$=C_{\mathcal{I}}\int_{\mathbb{R}_{+}}\widetilde{\mathbf{E}}_{x,i}\left[\frac{R_{J_{\xi}(s)}(b+X_{\xi}(s))}{R_{i}(b+x)}R_{J_{\xi}(s)}(b+X_{\xi}(s))\mathbbm{1}_{\left\{\inf_{r\leq s}X_{\xi}(r)\geq-b\right\}}V_{J_{\xi}(s)}(\theta^{*})e^{-\frac{2}{3}X_{\xi}(s)}\right]\mathrm{d}s.$
Since $\int R_{j}^{2}(b+x)V_{j}(\theta^{*})e^{-x}e^{x/3}\mathrm{d}x\leq C\int x^{2}e^{-2x/3}\mathrm{d}x<\infty$, by Lemma 2.2, $\widetilde{\mathbf{E}}^{\uparrow}_{x,i}[A_{1}]<\infty$, and therefore, $A_{1}<\infty$, $\widetilde{\mathbf{P}}^{\uparrow}_{x,i}$-a.s.
@@
$\widetilde{\mathbf{E}}^{\uparrow}_{x,i}[I]$ $\leq C_{\mathcal{I}}\widetilde{\mathbf{E}}^{\uparrow}_{x,i}\bigg[\int_{\mathbb{R}_{+}}\sum_{k\geq 0}\widetilde{\mu}_{J_{s}}(k)\mathbbm{1}_{\left\{k>e^{\frac{1}{3}X_{\xi}(s)}\right\}}\mathrm{d}s\bigg]$
$\leq C_{\mathcal{I}}\int_{\mathbb{R}_{+}}\widetilde{\mathbf{E}}^{\uparrow}_{x,i}\bigg[\sum_{k\geq 0}\widetilde{\mu}_{J_{s}}(k)\mathbbm{1}_{\left\{k>e^{\frac{1}{3}X_{\xi}(s)}\right\}}\bigg]\mathrm{d}s$
$=C_{\mathcal{I}}\int_{\mathbb{R}_{+}}\widetilde{\mathbf{E}}_{x,i}\bigg[\frac{R_{J_{\xi}(s)}(b+X_{\xi}(s))}{R_{i}(b+x)}\mathbbm{1}_{\left\{\inf_{r\leq s}X_{\xi}(r)\geq-b\right\}}\sum_{k\geq 0}\widetilde{\mu}_{J_{s}}(k)\mathbbm{1}_{\left\{k>e^{\frac{1}{3}X_{\xi}(s)}\right\}}\bigg]\mathrm{d}s.$
Applying Lemma 2.2 with $f(r,j):=(b+r)\sum_{k\geq 1}\widetilde{\mu}_{j}(k)\mathbbm{1}_{\left\{3\log k>r\right\}}$, we obtain
$\int_{-b}^{\infty}f(r,j)\mathrm{d}r$ $=\int_{-b}^{\infty}(b+r)\sum_{k\geq 1}\widetilde{\mu}_{j}(k)\mathbbm{1}_{\left\{3\log k>r\right\}}\mathrm{d}r$
$\leq\frac{1}{2}\sum_{k\geq 1}\max_{i\in\mathcal{I}}\widetilde{\mu}_{i}(k)(3\log k+b)^{2}.$
@@
$\sum_{k\geq 1}\max_{i\in\mathcal{I}}\widetilde{\mu}_{i}(k)(\log k)^{2}\leq\sum_{i\in\mathcal{I}}\sum_{k\geq 1}\widetilde{\mu}_{i}(k)(\log k)^{2}=\sum_{i\in\mathcal{I}}\frac{1}{m_{i}}\sum_{k\geq 1}{\mu}_{i}(k)k(\log k)^{2}<\infty.$
@@
Fix $\varepsilon>0$. By (3.10), there exists $b=b(\varepsilon)\geq 0$ such that $\mathbf{P}_{x,i}(\inf_{t\geq 0}\inf_{u\in\mathcal{N}_{t}}X_{u}(t)\geq-b\mid\mathscr{S})\geq 1-\varepsilon$.
@@
$Z_{\theta^{*}}(t)-Z^{\prime}_{\theta^{*}}(t)=-\sum_{u\in\mathcal{N}_{t}}e^{-X_{u}(t)}V’_{J_{u}(t)}(\theta^{*}).$ (3.12)
@@
$\mathbf{P}_{x,i}\left(Z_{\theta^{*}}(\infty)=0\right)=\mathbf{P}_{x,i}\left(\lim_{t\to\infty}\sum_{u\in\mathcal{N}_{t}}(X_{u}(t)-x)e^{-(X_{u}(t)-x)}V_{J_{u}(t)}(\theta^{*})=0\right)=\mathbf{P}_{0,i}(Z_{\theta^{*}}(\infty)=0).$
@@ 4  FKPP equations and travelling waves
$\Phi(x,i)=\mathbb{E}_{x+\rho t,i}\left[\Phi(\chi_{i}(t),i)\right]$
$+\int_{0}^{t}\mathbb{E}_{x+\rho t,i}\bigg[\beta_{i}g_{i}(\Phi(\chi_{i}(t-s)-\rho s,i))+\sum_{j\neq i}q_{ij}\Phi(U_{ij}+\chi_{i}(t-s)-\rho s,j)-(q_{i}+\beta_{i})\Phi(\chi_{i}(t-s)-\rho s,i)\bigg]\mathrm{d}s.$
@@
$\Phi(x,i)=$ $\,\mathbb{E}_{x,i}\left[\Phi(\widehat{\chi}_{i}(t),i)\right]$ (4.1)
$+\int_{0}^{t}\mathbb{E}_{x,i}\bigg[\beta_{i}g_{i}(\Phi(\widehat{\chi}_{i}(t\!-\!s),i))+\sum_{j\neq i}q_{ij}\Phi(U_{ij}+\widehat{\chi}_{i}(t\!-\!s),j)-(q_{i}+\beta_{i})\Phi(\widehat{\chi}_{i}(t\!-\!s),i)\bigg]\mathrm{d}s.$ (4.2)
@@
$0=\left({\cal A}_{i}+\rho\frac{\partial}{\partial x}\right)\Phi(x,i)+\int_{\mathbb{R}}\sum_{j\neq i}q_{ij}\big(\Phi(x+y,j)-\Phi(x,i)\big)\mathbb{P}(U_{ij}\in\mathrm{d}y)+\beta_{i}\big(g_{i}(\Phi(x,i))-\Phi(x,i)\big).$

### v2

@@ 3.1 Truncated derivative martingales
For $x\in\mathbb{R}$, let $\tau_{x}:=\inf\{t\geq 0\colon\widehat{X}_{\xi}(t)<x\}$. According to [19, Theorem 29], there exists finite and non-negative functions $(R_{i})_{i\in\mathcal{I}}$, such that the process
@@
Since $\widetilde{\mathbf{E}}_{x,i}^{\theta^{*}}[(-\widehat{X}_{\xi})^{2}(t)]<\infty$, we know from [19, Theorem 35 and Lemma 38] that $(-\widehat{X}_{\xi},J_{\xi})$ has tight overshoot under $\widetilde{\mathbf{P}}_{x,i}^{\theta^{*}}$, and equivalently, $\sum_{j\in\mathcal{I}}\widetilde{\pi}_{j}\widetilde{\mathbf{E}}_{0,i}^{\theta^{*}}[H^{-}(1)]\in(0,\infty)$, where $H^{-}$ is the so-called descending ladder height process of $(\widehat{X}_{\xi},J_{\xi})$ under $\widetilde{\mathbf{P}}_{0,i}^{\theta^{*}}$ and $\widetilde{\pi}$ is the stationary distribution of $J_{\xi}(t)$ under $\widetilde{\mathbf{P}}_{0,i}^{\theta^{*}}$.  [^6] Note that, as remarked by [19, Page 1995], we do not require non-lattice assumption. Then it follows from the Markov renewal theory (see e.g. [39] or [19, Theorem 28 (i)]; note that the version we use here holds for the lattice case as well) that
@@
$\mathbf{E}_{x,i}\left[Z_{\theta^{*}}^{(b)}(t)\right]$ $=V_{i}(\theta^{*})e^{-\theta^{*}x}\widetilde{\mathbf{E}}_{x,i}^{\theta^{*}}\left[R_{J_{\xi}(t)}(\widehat{X}_{\xi}(t)+b)\mathbbm{1}_{\left\{\inf_{s\leq t}\widehat{X}_{\xi}(s)\geq-b\right\}}\right]$
$=V_{i}(\theta^{*})e^{-\theta^{*}x}R_{i}(\theta^{*}x+b).$
@@
The convergence of the derivative martingale, stated in Theorem 1.6, would rely on the the following $\mathcal{L}^{1}$-convergence result of the truncated derivative martingale $Z_{\theta^{*}}^{(b)}$.
@@
Then $\mathbb{P}_{x,i}^{\uparrow}$ is known as the law of a MAP conditioned to stay positive; we refer to [19, Appendices A.7-A.8] for more details. Lévy processes (with single type) conditioned to stay positive have been the subject of a large literature; we refer to [14] and reference therein. Define $\mathbf{P}_{x,i}^{\uparrow}$ as the projection of $\widetilde{\mathbf{P}}_{x,i}^{\uparrow}$ on $(\mathcal{T},\mathcal{F})$. By projection (2.29) and change of measure (2.23), we also deduce the connection with the original law $\mathbf{P}_{x,i}$: for $t\geq 0$,
@@
Denote by $\mathbf{M}(\mathrm{d}s,\mathrm{d}k)$ the counting measure (on $\mathbb{R}_{+}\times\mathbb{Z}_{>0}$) of the spine’s fission times and number of children. Let $\mathcal{G}$ be the $\sigma$-field generated by the spine $(X_{\xi},J_{\xi})$ and $\mathbf{M}$, as given in Lemma 2.50. By [54, Lemma 4.2], to show that $(Z_{\theta^{*}}^{(b)}(t),t\geq 0)$ is uniformly integrable, and $\mathbf{E}_{x,i}\left[\frac{Z_{\theta^{*}}^{(b)}(\infty)}{Z_{\theta^{*}}^{(b)}(0)}\right]=1$, it suffices to prove that
@@
$\widetilde{\mathbf{E}}_{x,i}^{\uparrow}\left[Z_{\theta^{*}}^{(b)}(t)\,\middle|\,\mathcal{G}\right]=$ $\,R(b+X_{\xi}(t))e^{-X_{\xi}(t)}$
$+\int_{[0,t]\times\mathbb{Z}_{>0}}kR_{J_{\xi}(s)}(b+X_{\xi}(s))V_{J_{\xi}(s)}(\theta^{*})e^{-X_{\xi}(s)}\mathbf{M}(\mathrm{d}s,\mathrm{d}k).$
As $X_{\xi}(t)$ under $\widetilde{\mathbf{P}}^{\uparrow}_{x,i}$ is a MAP conditioned to stay above b-b, by [19, Proposition 33] we have $\lim_{t\to\infty}X_{\xi}(t)=+\infty$ $\widetilde{\mathbf{P}}^{\uparrow}_{x,i}$-a.s. So the first term above converges to 0, $\widetilde{\mathbf{P}}^{\uparrow}_{x,i}$-a.s.
@@
Recall that $\mathbf{M}$ is a Cox process, in the sense that given $(X_{\xi}(t),{J}_{\xi}(t))_{t\geq 0}$, the conditional distribution of $\mathbf{M}$ is a Poisson point process with intensity $\widetilde{\beta}_{{J}_{\xi}(t)}\mathrm{d}t\otimes\mathrm{d}\widetilde{\mu}_{{J}_{\xi}(t)}$, with each $\widetilde{\mu}_{i}$ the size-biased offspring distribution. Let $C_{\mathcal{I}}:=\max_{i\in\mathcal{I}}\widetilde{\beta}_{i}>0$. Then the intensity is bounded by $C_{\mathcal{I}}\mathrm{d}t\otimes\mathrm{d}\widetilde{\mu}_{{J}_{\xi}(t)}\mathbbm{1}_{\left\{\widetilde{\beta}_{J_{\xi}(t)}>0\right\}}$. The compensation formula leads to
$\hskip 15.00002pt\widetilde{\mathbf{E}}^{\uparrow}_{x,i}[A_{1}]$
$\leq C_{\mathcal{I}}\widetilde{\mathbf{E}}^{\uparrow}_{x,i}\bigg[\int_{\mathbb{R}_{+}}\!\!R_{J_{\xi}(s)}(b\!+\!X_{\xi}(s))V_{J_{\xi}(s)}(\theta^{*})e^{-X_{\xi}(s)}\Big(\sum_{k\geq 1}\widetilde{\mu}_{J_{s}}(k)k\mathbbm{1}_{\left\{k\leq e^{\frac{1}{3}X_{\xi}(s)}\right\}}\Big)\mathbbm{1}_{\left\{\widetilde{\beta}_{J_{\xi}(s)}>0\right\}}\mathrm{d}s\bigg]$
$\leq C_{\mathcal{I}}\int_{\mathbb{R}_{+}}\widetilde{\mathbf{E}}^{\uparrow}_{x,i}\left[R_{J_{\xi}(s)}(b+X_{\xi}(s))V_{J_{\xi}(s)}(\theta^{*})e^{-\frac{2}{3}X_{\xi}(s)}\mathbbm{1}_{\left\{\widetilde{\beta}_{J_{\xi}(s)}>0\right\}}\right]\mathrm{d}s$
$=C_{\mathcal{I}}\int_{\mathbb{R}_{+}}\widetilde{\mathbf{E}}_{x,i}\left[\frac{R_{J_{\xi}(s)}(b+X_{\xi}(s))}{R_{i}(b+x)}R_{J_{\xi}(s)}(b+X_{\xi}(s))\mathbbm{1}_{\left\{\inf_{r\leq s}X_{\xi}(r)\geq-b\right\}}V_{J_{\xi}(s)}(\theta^{*})\right$.
$\hskip 80.00012pt\left.e^{-\frac{2}{3}X_{\xi}(s)}\mathbbm{1}_{\left\{\widetilde{\beta}_{J_{\xi}(s)}>0\right\}}\right]\mathrm{d}s.$
Since $\int R_{j}^{2}(b+x)V_{j}(\theta^{*})e^{-x}e^{x/3}\mathrm{d}x\leq C\int x^{2}e^{-2x/3}\mathrm{d}x<\infty$, it follows from Lemma 2.2 that $\widetilde{\mathbf{E}}^{\uparrow}_{x,i}[A_{1}]<\infty$. Therefore, we have $A_{1}<\infty$, $\widetilde{\mathbf{P}}^{\uparrow}_{x,i}$-a.s.
@@
$\widetilde{\mathbf{E}}^{\uparrow}_{x,i}[I]\leq C_{\mathcal{I}}\widetilde{\mathbf{E}}^{\uparrow}_{x,i}\bigg[\int_{\mathbb{R}_{+}}\sum_{k\geq 0}\widetilde{\mu}_{J_{s}}(k)\mathbbm{1}_{\left\{k>e^{\frac{1}{3}X_{\xi}(s)}\right\}}\mathbbm{1}_{\left\{\widetilde{\beta}_{J_{\xi}(s)}>0\right\}}\mathrm{d}s\bigg]$
$\leq C_{\mathcal{I}}\int_{\mathbb{R}_{+}}\widetilde{\mathbf{E}}^{\uparrow}_{x,i}\bigg[\sum_{k\geq 0}\widetilde{\mu}_{J_{s}}(k)\mathbbm{1}_{\left\{k>e^{\frac{1}{3}X_{\xi}(s)}\right\}}\mathbbm{1}_{\left\{\widetilde{\beta}_{J_{\xi}(s)}>0\right\}}\bigg]\mathrm{d}s$
$=C_{\mathcal{I}}\int_{\mathbb{R}_{+}}\widetilde{\mathbf{E}}_{x,i}\bigg[\frac{R_{J_{\xi}(s)}(b+X_{\xi}(s))}{R_{i}(b+x)}\mathbbm{1}_{\left\{\inf_{r\leq s}X_{\xi}(r)\geq-b\right\}}\sum_{k\geq 0}\widetilde{\mu}_{J_{s}}(k)\mathbbm{1}_{\left\{k>e^{\frac{1}{3}X_{\xi}(s)}\right\}}\mathbbm{1}_{\left\{\widetilde{\beta}_{J_{\xi}(s)}>0\right\}}\bigg]\mathrm{d}s.$
Applying Lemma 2.2 with $f(r,j):=(b+r)\sum_{k\geq 1}\widetilde{\mu}_{j}(k)\mathbbm{1}_{\left\{3\log k>r\right\}}\mathbbm{1}_{\left\{\widetilde{\beta}_{j}>0\right\}}$, we obtain
$\int_{-b}^{\infty}f(r,j)\mathrm{d}r$ $=\int_{-b}^{\infty}(b+r)\sum_{k\geq 1}\widetilde{\mu}_{j}(k)\mathbbm{1}_{\left\{3\log k>r\right\}}\mathbbm{1}_{\left\{\widetilde{\beta}_{j}>0\right\}}\mathrm{d}r$
$\leq\frac{1}{2}\sum_{k\geq 1}\max_{i\in\mathcal{I}}\left\{\widetilde{\mu}_{i}(k)\mathbbm{1}_{\left\{\widetilde{\beta}_{i}>0\right\}}\right\}(3\log k+b)^{2}.$
@@
$\sum_{k\geq 1}\max_{i\in\mathcal{I}}\left\{\widetilde{\mu}_{i}(k)\mathbbm{1}_{\left\{\widetilde{\beta}_{i}>0\right\}}\right\}(\log k)^{2}$ $\leq\sum_{i\in\mathcal{I}}\sum_{k\geq 1}\widetilde{\mu}_{i}(k)\mathbbm{1}_{\left\{\widetilde{\beta}_{i}>0\right\}}(\log k)^{2}$ (3.10)
$=\sum_{i\in\mathcal{I}}\frac{1}{m_{i}}\mathbbm{1}_{\left\{\widetilde{\beta}_{i}>0\right\}}\sum_{k\geq 1}{\mu}_{i}(k)k(\log k)^{2}<\infty.$ (3.11)
@@
Fix $\varepsilon>0$. By (3.12), there exists $b=b(\varepsilon)\geq 0$ such that
$\mathbf{P}_{x,i}(\inf_{t\geq 0}\inf_{u\in\mathcal{N}_{t}}X_{u}(t)\geq-b\mid\mathscr{S})\geq 1-\varepsilon.$
@@
$Z_{\theta^{*}}(t)-Z^{\prime}_{\theta^{*}}(t)=-\sum_{u\in\mathcal{N}_{t}}e^{-X_{u}(t)}V^{\prime}_{J_{u}(t)}(\theta^{*}).$ (3.14)
@@
$\mathbf{P}_{x,i}\left(Z_{\theta^{*}}(\infty)=0\right)$ $=\mathbf{P}_{x,i}\left(\lim_{t\to\infty}\sum_{u\in\mathcal{N}_{t}}(X_{u}(t)-x)e^{-(X_{u}(t)-x)}V_{J_{u}(t)}(\theta^{*})=0\right)$ (3.15)
$=\mathbf{P}_{0,i}(Z_{\theta^{*}}(\infty)=0).$ (3.16)
@@ 4  FKPP equations and travelling waves
$\Phi(x,i)=\mathbb{E}_{x+\rho t,i}\left[\Phi(\chi_{i}(t),i)\right]+\int_{0}^{t}\mathbb{E}_{x+\rho t,i}\bigg[\beta_{i}g_{i}(\Phi(\chi_{i}(t-s)-\rho s,i))$
$\hskip 60.00009pt+\sum_{j\neq i}q_{ij}\Phi(U_{ij}+\chi_{i}(t-s)-\rho s,j)-(q_{i}+\beta_{i})\Phi(\chi_{i}(t-s)-\rho s,i)\bigg]\mathrm{d}s.$
@@
$\Phi(x,i)=$ $\,\mathbb{E}_{x,i}\left[\Phi(\widehat{\chi}_{i}(t),i)\right]+\int_{0}^{t}\mathbb{E}_{x,i}\bigg[\beta_{i}g_{i}(\Phi(\widehat{\chi}_{i}(t\!-\!s),i))$ (4.1)
$\qquad+\sum_{j\neq i}q_{ij}\Phi(U_{ij}+\widehat{\chi}_{i}(t\!-\!s),j)-(q_{i}+\beta_{i})\Phi(\widehat{\chi}_{i}(t\!-\!s),i)\bigg]\mathrm{d}s.$ (4.2)
@@
$0=$ $\,\left({\cal A}_{i}+\rho\frac{\partial}{\partial x}\right)\Phi(x,i)+\int_{\mathbb{R}}\sum_{j\neq i}q_{ij}\big(\Phi(x+y,j)-\Phi(x,i)\big)\mathbb{P}(U_{ij}\in\mathrm{d}y)$
$\,+\beta_{i}\big(g_{i}(\Phi(x,i))-\Phi(x,i)\big).$